Calico prints custom training material for corporations. The business was started January 1, 2017. The company uses a normal costing system. It has two direct cost pools, materials and labor, and one indirect cost pool, overhead. Overhead is charged to printing jobs on the basis of direct labor cost. The following shows the information available for 2017.

Budgeted direct labor costs              $180,000
Budgeted overhead costs                  $324,000
Costs of actual material used           $155,000
Actual direct labor costs                    $171,000
Actual overhead costs                        $308,500

There were two jobs in process on December 31, 2017: Job 11 and Job 12. Costs added to each job as of December 31 are as follows:

                     Direct materials                     Direct Labor
Job 11       $4,900                                       $4,800
Job 12       $5,630                                       $6,600

1. Compute the overhead allocation rate.
2. Calculate the balance in ending work in process and cost of goods sold before any adjustments for under or overallocated overhead.
3. Calculate under or overallocated overhead
4. Calculate the ending balances in work in process and cost if goods sold if the under or overallocated amount is as follows:
    a. Written off to cost if goods sold.
    b. Prorated using the overhead allocated in 2017 (before proration) in the ending balances if cost of goods sold and work in process control accounts.
5. Which of the methods in requirement 4 would you choose? Explain
Answer #1

Part 1

overhead allocation rate = Budgeted overhead costs / Budgeted direct labor costs= 324000/180000 = 180% of direct labor cost

Part 2

Job 11

Job 12

Total

Direct material costs

4900

5630

10530

Direct labor costs

4800

6600

11400

Overhead (180%* direct labor cost)

8640

11880

20520

Total costs

18340

24110

42450

Cost of goods sold = Beginning WIP + Manufacturing costs – Ending WIP = 0+(155000+171000+(171000*180%))-42450 = $591350

Part 3

Overhead allocated = 171000*180% = $307800

Underallocated overhead = Actual overhead – Allocated overhead= $308500 – $307800 = $700 underallocated

Part 4 a

Work in Process

$42450

Cost of goods sold

$592050

Account(1)

Dec. 31, 2014 Account Balance (Before Proration)(2)

Write-off of $700

Underallocated overhead(3)

Dec. 31, 2014Account Balance(After Proration)(4) = (2) + (3)

Work in Process

42450

0

42450

Cost of goods sold

591350

700

592050

633800

700

634500

Part 4 b

Work in Process

$42499

Cost of goods sold

$592001

Account(1)

Dec. 31, 2014 Account Balance (Before Proration)

Allocated Overhead Included in Dec. 31, 2014 Account Balance(Before Proration)

Write-off of $700

Underallocated overhead

Dec. 31, 2014Account Balance(After Proration)

(1)

(2)

(3)

(4)

(5)

(6) = (2) + (5)

Work in Process

42450

20520

7.04%

49

42499

Cost of goods sold

591350

270864

92.96%

651

592001

633800

291384

100.00%

700

634500

(171000-20520)*180% = 270864

Part 5

Calico should write off the $700 underallocated overhead to Cost of Goods Sold account. When cost of goods sold is large relative to Work-in-Process Inventory and Finished Goods Inventory and also if the underallocated overhead is immaterial, then it is better to write off all of the underallocated overhead to Cost of Goods Sold. In the given case both conditions are applicable.